Shift summary for owl 10/8/99

Philippe Gouffon (pgouffon@fdei01)
Fri, 08 Oct 1999 07:11:16 -0500

Shift Summary for Owl Shift, 10/8/99

On Shift: Philippe Gouffon, Bob Tschirhart

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

========================================================================
Overview of shift:
No beam. TeV purification in progress.

========================================================================
Description of Shift:

- Contribution of Brazil: 1kg of coffee beans delivered to Ktev Hall. Some
of it tested.
- Installed back the EDM battery and measured the distances
- Examined pulse height and timing on accidental counters at station #1.
Threshold for discriminator is 30mV, a litlle low but well above noise.
- around 3AM, CSi block 1533 gain started to fluctuate. At 0445 it dropped
out of the laser illumination plot. Julie notified and the voice silenced.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

1) Fix channel 1533

2) Ar Ethane (West) pressure is ZERO since yesterday (10/7) evening. May
be bottles should be changed.

- ------- End of Forwarded Message

------- End of Forwarded Message